OpenStudy (valirium):

Find all the zeroes of the equation. -3x^4+27x^2+1200=0

OpenStudy (valirium):

@Vuriffy

OpenStudy (518nad):

y=x^2

OpenStudy (518nad):

-3y^2+27y+1200=0 -3(y^2-9y-400)=0 y^2-9y-400=0 can u factor this, and find solution for y

OpenStudy (valirium):

(y-25)(y+16)=0

OpenStudy (alyssa1999xo):

and since y=x^2 \((x^2-25)(x^2+16)=0\) from there, we can set each term equal to 0 and solve \(x^2-25=0\\x^2=25\\x=\sqrt25\\x=-5,5\) \((x^2+16)=0 \\x^2=-16 \\x=\sqrt{-16} \\x=4i,-4i \)

OpenStudy (mathmale):

Note that the given equation, -3x^4+27x^2+1200=0, can be factored by dividing each of the four terms by -3. Try doing this. I agree it's a good approach to let x^2 = y. Having factored out -3, and subst. y for x^2, what do you get?

OpenStudy (valirium):

\(-3(y^2)- 9y- 400)= 0 \) \(y^2)- 9y- 400= 0 \) @mathmale